Zinc metal reacts with hydrochloric acid according to the following balanced equation.

Zn(s) + 2HCl(aq) --> ZnCl_2(aq) + H_2(g)
When 0.124 g of Zn(s) is combined with enough HCl to make 51.3 mL of solution in a coffee-cup calorimeter, all of the zinc reacts, raising the temperature of the solution from 21.5 C to 23.9 C.
Find Delta Hrxn for this reaction as written. (Use 1.0 g/mL for the density of the solution and 4.18 J/g *C as the specific heat capacity.)
Answer in kJ/mol
Answer #1

Delta Hrxn = total heat given out/ number of moles of zn.
moles of zn = 0.124/65 = 1.908*10-3 mol

heat given out = mass * change in temp * specific heat capacity

                       = (51.3*1.0) * (23.9 - 21.5) * 4.18 = 514.6 J

Delta Hrxn = - (514.6/1.908*10-3 ) = - 269.7 kJ/mol

negative cos heat is given out.